OpenStudy (anonymous):

f(x) ln x+2; g(x)= e^x-2 How to show how f and g are inverses of each other?

OpenStudy (anonymous):

compose them and see that you get the identity back

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(e^{x-2})=\ln(e^{x-2})+2=x-2+2=x\] in one direction

OpenStudy (anonymous):

i hope that is clear, i wrote 5 equal signs first one is definition of the circle notations second one is replacing the generic \(f\) by the specific one you have third is evaluating \(f\) at \(e^{x-2}\) fourth is because log and exp are inverse functions ' fifth is arithmetic

OpenStudy (anonymous):

Got it, Thank you!

OpenStudy (anonymous):

yw i made a typo in my response however, should have been "second one is replacing he generic \(g\) by the specific one "